# Reason3D: Searching and Reasoning 3D Segmentation via Large Language Model

### v1

@@ B.2 Data Statistics
B.5 More Visualization Results.
3D Reasoning Segmentation. We provide more qualitative examples for the 3D reasoning segmentation task and the predictions by our Reason3D in Figure 5.
@@
Figure 7: Failure cases. (a.) Small queried objects. (b.) Similar object in the scene. (c.) Similar structures of the point cloud. (d.) The question requiring complicate world knowledge. The purple regions denotes the predicted segmentation masks from our Reason3D and the green means the ground truth. Best viewed with zoom in.
Appendix D Limitations
While our model introduces a novel approach to the 3D reasoning segmentation task, it does have several limitations. Firstly, we observe that our model cannot handle very large-scale scenes for 3D reasoning segmentation (e.g., searching for an object within a house with 30 rooms in the Matterport dataset). Furthermore, our model lacks the ability to handle false premises cases (e.g., searching for an object that may not be inside the scenes), which could be a promising direction to explore within the LLM-based 3D reasoning framework.
Appendix E Broader Impact
Reason3D is designed to segment objects in 3D space based on language inputs. Compared to traditional 3D segmentation algorithms, Reason3D models have a lower customization threshold, allowing users to identify objects through natural language. However, this accessibility could potentially be abused. Additionally, the datasets and pre-trained models utilized in Reason3D may contain inherent biases, which could be reflected in the model’s performance.

### v2

@@ B.4 Instruction Template
3D Reasoning Segmentation: “Human: <scene> can you segment the object in the scene with the following descriptions: description? Assistant: Sure, it’s <SEG>.”
3D Hierarchical Searching. “Human: <scene> can you segment the object based on the description: description? Please segment the target room first, then output the object mask. Assistant: Sure, the room is <LOC>, and the object is <SEG>.”
3D Express Referring Segmentation. “Human: <scene> please segment the object from the given scene: description. Assistant: It’s <SEG>."
3D Question Answering. “Human: <scene> please answer the question based on the given scene: question and output the related segmentation mask. Assistant: answer <SEG>."
B.5 3D Question Answering Results
Evaluation Metrics. For the QA task, the evaluation metrics include BLEU-4 [34], ROUGE-L [30], METEOR [5], and CIDEr [50] to ensure robust answer matching. These metrics evaluate the precision, fluency, and semantic accuracy of the responses.
Results. In addition to excelling in 3D reasoning and referring tasks, our approach also performs well in 3D question answering tasks. We present our results on the ScanQA validation set in Table 9, where we observed a significant improvement in evaluation metrics over both baseline methods and the recent LLM-based method, 3D-LLM [19]. Our approach not only answers questions accurately but also visualizes the related segmentation masks to further demonstrate the effectiveness.
B.6 3D Visual Grounding Results.
Evaluation Metrics. Similar to the 3D expressive referring task, the evaluation metrics used are Acc@0.25 and Acc@0.5, indicating the percentage of correctly predicted bounding boxes with an IoU greater than 0.25 or 0.5, respectively, compared to the ground truth.
Results. Although our primary focus is on 3D segmentation, our method also effectively predicts 3D bounding boxes as supplementary outputs, facilitating comparison with 3D visual grounding methods. To generate a 3D bounding box for a referred object, we first apply DBSCAN [15] to eliminate noisy points, and then calculate the minimum and maximum XYZ coordinates from the points within the segmentation mask to form the 3D box. As demonstrated in Table 9, our approach not only outperforms recent 3D visual grounding methods but also significantly surpasses LLM-based methods, such as 3D-LLM [19], which struggle to integrate textual and numerical data to accurately localize objects in 3D space.
B.7 More Visualization Results.
3D Reasoning Segmentation. We provide more qualitative examples for the 3D reasoning segmentation task and the predictions by our Reason3D in Figures 5 and 6.
@@
Figure 6: More Visualization Results for 3D Reasoning Segmentation Tasks. The purple regions highlight the predicted segmentation masks generated by our model. Best viewed with zoom in.
@@
Figure 8: Failure cases. (a.) Small queried objects. (b.) Similar object in the scene. (c.) Similar structures of the point cloud. (d.) The question requires complicated world knowledge. The purple regions denote the predicted segmentation masks from our Reason3D, and the green means the ground truth. Best viewed with zoom in.
D Limitations
While our model introduces a novel approach to 3D reasoning segmentation, it does have limitations that open promising avenues for future research. For example, our system currently struggles with large-scale scenes—such as identifying an object within a 30-room house in the Matterport dataset. Future work could refine the hierarchical decoder or integrate adaptive multi-scale processing techniques to better manage complex spatial extents. Additionally, the model does not yet handle scenarios with false premises—such as querying for an object that may not be present—suggesting that incorporating uncertainty estimation or robust error-detection mechanisms could help validate query assumptions before processing. Moreover, since our implementation is designed primarily for single-object queries, its performance on multi-object or multi-category tasks remains untested. This limitation points to the need for developing enriched query representations and joint optimization strategies that can simultaneously manage multiple objects. Addressing these challenges would significantly enhance the robustness and versatility of our LLM-based 3D reasoning framework.
E Broader Impact
Reason3D is designed to segment objects in 3D space based on language inputs. Compared to traditional 3D segmentation algorithms, Reason3D models have a lower barrier to customization, enabling users to identify objects using natural language. However, this increased accessibility also raises the potential for misuse. Furthermore, the datasets and pre-trained models used in Reason3D may carry inherent biases, which could influence the model’s performance.